Please help with the following question from a quiz review:
The function d=125+80n models the amount of money, in dollars, the zip line ticket cashier has in her cash register after selling n tickets. How much money did the cashier have in the register before she sold any tickets? Please explain and show as much work as possible!
Thank you!!

Answer by 31(5)

In this problem, "d" equals the total amount of money the zip line cashier has in her register and "n" represents the number of tickets she sold. Therefore, if you want to know how much money she had before the ticket sales, you would substitute n=O into the equation (since you want to know how much she had when she sold zero tickets).

d=125+80n
d=125+80(0)
d=125+0
d=125

Final Answer: The cashier had $125 in her register before selling any tickets.

Hope this helped!